# Minimax discrete distribution estimation with self-consumption

### v1

@@ IV-A The error term $h_{t}({\nicefrac{{1}}{{4k}}})$
The multi-stage lower bounds depend on the performance of estimators in previous stages through the probabilities $g_{i-1}\left(\nicefrac{{1}}{{4k}}\right)$. It is perhaps surprising that the lower bounds decrease as these probabilities increase. The source of this apparent improvement is the looser upper bound on the KL divergence in (11) when the previous estimators perform badly. This suggests that it might be easier for an estimator to perform well when the synthetic samples are from a distribution sufficiently removed from the real one.
The estimator analyzed in the upper bound section does not share this behavior, leading to a gap between the upper and lower bounds in the case where the proportion of real samples decays very rapidly. The estimator is also kept unbiased at every stage via assumptions on the problem parameters; this helps keep the analysis from becoming almost prohibitively cumbersome. A careful analysis of a more general class of linear estimators might help bridge the gap with the lower bounds or drop the assumptions required. Alternatively, a tighter analysis bounding the aforementioned divergence might lead to a lower bound that matches the upper bound wherever valid.
IV-B Comparison against oracle-assisted minimax loss
To understand the impact of not knowing the source identity on the minimax loss, it is useful to compare the upper and lower bounds presented above to the ones when the identity of the source of each sample is known in the form of an auxiliary random vector $W:=W_{i}^{n_{i}}\sim(\mathrm{Bern}(\alpha_{i}))^{n_{i}},i\in\{1,\dots,t\}$ provided to the estimator by an oracle.
Lemma 4 (Oracle-assisted minimax loss)
There exist positive constants $c_{1},c_{2}$ such that the minimax loss $r^{\ell_{2}^{2},\mathrm{oracle}}_{t,k}$ of self-consuming estimation with information on source identity is bounded between
$\frac{c_{1}}{\sum_{i=0}^{t}n_{i}\alpha_{i}}\leq r^{\ell_{2}^{2},\mathrm{oracle}}_{t,k}\leq\frac{c_{2}}{\sum_{i=0}^{t}n_{i}\alpha_{i}}.$
@@ IV-B Comparison against oracle-assisted minimax loss
Using Lemma 5 and the chain rule of the KL divergence, along with the fact that
$D(\tilde{P}({p_{v},\hat{P}^{(i)},\alpha})\|\tilde{P}({p_{v-2e_{j}},\hat{P}^{(i)},\alpha})|W)=\alpha D(p_{v}\|p_{v-2e_{j}})$
leads to an $\ell_{2}^{2}$ lower bound of $\frac{1}{512\sum_{i=0}^{t}n_{i}\alpha_{i}}$. The empirical distribution of the real samples can be shown to have an $\ell_{2}^{2}$ loss within a constant factor of $\frac{1-\nicefrac{{1}}{{k}}}{\sum_{i=0}^{t}n_{i}\alpha_{i}}$ via a Chernoff-style argument. Similar bounds can be obtained for the $\ell_{1}$ loss. ∎
The absence of source information at the estimator leads to a penalty factor of roughly $\max(\alpha_{i},g_{i-1}(\nicefrac{{1}}{{4k}}))$ in the ‘effective sample size’ of each batch. This can lead to large significant gaps between the base and oracle-assisted losses. Depending on whether or not the $\alpha_{i}$’s are comparable to the $g_{i-1}(\nicefrac{{1}}{{4k}})$’s, there might also be a gap between the guarantees provided by the upper and the lower bounds. We consider a few examples of such cases in the sequel.
IV-C Examples
We demonstrate a few examples where the bounds in the main result of this paper are useful. Since the bounds for $\ell_{1}$ loss are $\sqrt{k}$ times the square root of the $l_{2}^{2}$ bounds, we only consider $\ell_{2}^{2}$ bounds here.
IV-C1 Data accumulation
We first demonstrate choices for the sequence $n_{t},\alpha_{t},t\geq 0$ where the upper and lower bounds match (up to constant factors) and show a significant gap from the oracle-assisted minimax loss in the accumulation setting. We also note how modifying these choices leads to a gap between the upper and lower bounds for the base case.
Matching bounds
Let $n_{i}=a(i+1)^{\beta+\gamma}$, and $\alpha_{i}=(i+1)^{-\gamma}$, where $a,\beta,\gamma$ are positive constants such that $a\geq\frac{k}{4}$ and $\gamma\leq\min(1,\nicefrac{{1{+}\beta}}{{2}})$. Thus, on an average, the number of synthetic samples is $\Theta(t^{\beta+\gamma})$, while the number of real samples is $\Theta(t^{\beta})$. The assumption on $a$ ensures that the lower bound (Theorem 1) holds for every $t\geq 1$, and the assumptions on $\gamma$ ensure that the upper bound (Theorem 2) is valid and the error term $h_{\nicefrac{{1}}{{4k}}}(t)$ is, in the worst case, comparable to $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}$ for each $t$. Additionally, for the base setting, assume that the estimators in Theorem 2 are used at each stage.
Using Lemma 4,
$\frac{c_{1}(1+\beta)}{a(t+2)^{1+\beta}}\leq r^{\ell_{2}^{2},\mathrm{oracle}}_{t,k}\leq\frac{c_{2}(1+\beta)}{a(t+1)^{1+\beta}}.$
In contrast, for the base case, using Theorem 2,
@@ IV-C1 Data accumulation
Chebyshev’s inequality leads to the following upper bound on the error term:
$h_{t}(\nicefrac{{1}}{{4k}})\leq{16k^{2}(\beta-\gamma+1)}\frac{t^{\gamma}}{\gamma}.$
@@
This inductively shows that the upper and lower bounds match in terms of order w.r.t. the stage $t$: for each stage $t$, the upper bound is within a constant factor of the lower bound, provided that order-optimal estimators are chosen for stages 00 through $t-1$. Moreover, there is a gap between the base and oracle-assisted minimax losses:
$\frac{r^{\ell_{2}^{2}}_{t,k}}{r^{\ell_{2}^{2},\mathrm{oracle}}_{t,k}}=\Theta(t^{\gamma})$ (7)
Bounds with a gap
@@
Consider the previous example but with $1>\gamma>\frac{1{+}\beta}{2}$ instead. The upper bound on $\gamma$ ensures that the upper bound of $\Theta((t+1)^{\beta+1-\gamma})$ from Theorem 2 is still valid, but the error term will now dominate the analysis of the lower bound (Theorem 1).
Let us assume that there exists a sequence of estimators for which the worst case loss is within a universal constant factor $d$ of the corresponding lower bound, and that these estimators are chosen at every stage. Let the worst-case expected loss of this sequence of estimators at stage $t$ be denoted by $\nicefrac{{\bar{\sigma}_{t}^{2}}}{{512}}$. Using Theorem 1, for each $i\geq 1$,
$\bar{\sigma}_{i}^{2}\geq\frac{1/a}{\frac{d}{\bar{\sigma}_{i-1}^{2}}+(i+1)^{\beta-\gamma}+\frac{(i+1)^{\beta}k^{2}\bar{\sigma}_{i-1}^{2}}{32}},$ (8)
and additionally, $\bar{\sigma}_{i}^{2}$ is at most a factor of $d$ times the RHS. This suggests $\bar{\sigma}_{i}^{2}$ is $\Theta\left((i+1)^{-s_{0}}\right)$ for some constant $s_{0}$. Proceeding inductively, assume (8) for $i=t-1$. Then $\nicefrac{{1}}{{\bar{\sigma}_{t-1}^{2}}}$ is $\Theta(t^{s_{0}})$, $(t+1)^{\beta-\gamma}$ is strictly smaller than $(t+1)^{\nicefrac{{\beta{-}1}}{{2}}}$ and $(t+1)^{\beta}\bar{\sigma}_{t-1}^{2}$ is $\Theta((t+1)^{\beta-s_{0}})$. Plugging in $s_{0}=\nicefrac{{\beta{+}1}}{{2}}$ satisfies the recursion, giving us a minimax loss of $\Theta(t^{\nicefrac{{\beta{+}1}}{{2}}})$.
Comparing this against the oracle-assisted minimax loss shows that there is a gap of at least $\Theta(t^{\nicefrac{{\beta{+}1}}{{2}}})$. The gap with the upper bound remains $\Theta(t^{\gamma})$. Since the error term $h_{1/4}(\cdot)$ dominates the analysis of the lower bound, the penalty for a much faster decaying sequence of $\alpha_{t}$’s might be limited. This is not to say that the impact of the penalty is small: if a target loss $L$ is achieved in $B_{0}$ batches in the oracle-assisted case, it will take the estimators matching the loss in Theorem 1 a number of batches in the order of at least $B_{0}^{2}$ to achieve the same fidelity. Proving that either of the bounds is tight, even in such extreme cases, remains an open problem.
We should also note that stronger upper bounds of the form $h_{t}(\nicefrac{{1}}{{4k}})\leq\exp(-\lambda_{0}t)$ can be found for some $\lambda_{0}>0$ if we assume that the estimators proposed in Theorem 2 are used in stages 00 through $t-1$. This would then make the same class of estimators order-optimal at stage $t$ for every $i\geq 0$. However, proving this relies on the fact that these estimates are subgaussian, and we would not be able to proceed if the estimates chosen in the previous stages have heavier than gaussian tails (but order-optimal error). The two regimes for $\beta,\gamma\in[0,1]^{2}$ are depicted in Figure 2.
Figure 2: Regimes of minimax loss with $\Theta(t^{\beta})$ real and $\Theta(t^{\beta+\gamma})$ synthetic samples at stage $t$; gaps indicated are w.r.t. the oracle-assisted loss
IV-C2 Data replacement
Consider the data replacement setting where, at every stage $t$, the estimator only has access to samples generated in the current stage $X_{t}^{n}$. Note that $\alpha$ and $n$ do not change with $t$ in this setting, except at stage 00 where all $n$ samples are real.
@@ IV-C2 Data replacement
Let the worst-case expected $\ell_{2}^{2}$ loss at stage $t$ be denoted as $\bar{\sigma}_{t}^{2}$, and define $M:=\sup_{t\geq 0}\bar{\sigma}_{t}^{2}$. Corollary 3, along with Chebyshev’s inequality then gives

### v2

@@ IV Discussion
It is evident from the bounds shown above that if the mixing factor $\alpha_{t}$ is constant (or even $\Omega(1)$), we can show that the base and oracle-assisted minimax losses are within constant factors of each other; the case where the probability converges to zero is more interesting. This corresponds to cases where the rate of synthetic sampling outpaces the real sampling rate by a considerable margin. Such cases might arise when it is significantly cheaper to sample the synthetic distribution, perhaps also exacerbated by the fact that the estimates get closer to the real distribution with every new batch. The behavior of the bounds in such cases depends on the rates of growth of $\sum_{i=0}^{t}n_{i}\alpha_{i}$ and $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}$. The absence of source information at the estimator leads to a penalty factor of roughly $\max(\alpha_{i},g_{i-1}(\nicefrac{{1}}{{4k}}))$ in the ‘effective sample size’ of each batch; we show that this can lead to large significant gaps between the base and oracle-assisted losses. Depending on whether or not the $\alpha_{i}$’s are comparable to the $g_{i-1}(\nicefrac{{1}}{{4k}})$’s, there might also be a gap between the guarantees provided by the upper and the lower bounds. In the sequel, we comment on the impact of the error term on the behavior of the bounds and the restrictions on the order-optimal estimator, show cases where differences in the aforementioned rates of growth imply gaps between the base and oracle-assisted losses, and conclude with a summary of the regimes where it is possible to comment precisely on the minimax error.
@@ IV-A The error term $h_{t}({\nicefrac{{1}}{{4k}}})$
The multi-stage lower bounds depend on the performance of estimators in previous stages through the probabilities $g_{i}\left(\nicefrac{{1}}{{4k}}\right)$. We note that the error term is small compared to the other term in the denominator, $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}$, in quite a lot of cases. Using Chebyshev’s inequality (as demonstrated in the upcoming examples) leads to an upper bound on each of these probabilities in terms of the minimax error $r_{i-1,k}^{\ell}$; the upper and lower bounds thus match whenever $n_{i}\alpha_{i}^{2}$’s are larger than or comparable to $1/\sum_{j=0}^{i-1}n_{j}\alpha_{j}^{2}$. We also note that it is necessary to use Chebyshev’s inequality—as opposed to stronger concentration results such as Hoeffding-style bounds that rely on the specific properties, such as subgaussianity in the case of the Hoeffding bound, of the estimators—to derive a universal upper bound on these probabilities for any sequence of estimators that achieves order-optimal performance.
@@
Intuitively, the quantities $g_{i}(\nicefrac{{1}}{{4k}})$ capture the worst-case deviation probability of any component of $\hat{P}_{i-1}$ from $p$ by more than $\nicefrac{{1}}{{4k}}$. Since we focus consider cases where $\hat{P}_{t}$ converges to $p$, this probability is expected to decay to zero. It is perhaps surprising, then, to also observe that the lower bounds decrease as the probabilities $g_{i}(\nicefrac{{1}}{{4k}})$ increase—suggesting that it might be possible to improve the performance of future estimators by estimating badly in the present. The source of this apparent improvement is the loosening of the upper bound on the KL divergence in (16) in instances when the previous estimators perform badly. This suggests that it might be easier for an estimator to perform well when the synthetic samples are from a distribution sufficiently removed from the real one; the sequence of estimators in Theorem 2 does not share this behavior.
IV-B The order-optimal estimator
The sequence of estimators in Theorem 2 do not always return distributions; indeed, an assumption on the sample sizes and mixing factors is required to make sure that the outputs stay in $\Delta_{k}$ at every stage. Each estimate in the sequence is also kept unbiased and uncorrelated from the previous estimate; this helps keep the analysis from becoming almost prohibitively cumbersome. A careful analysis of a more general class of nonlinear estimators might help bridge the gap with the lower bounds, proving the error term necessary, or drop the assumptions required. Alternatively, a tighter analysis bounding the aforementioned divergence might lead to a lower bound that matches the upper bound wherever valid. For the rest of this paper, we analyze the lower bounds assuming ‘good’ estimators—estimators with worst-case loss within at most a constant factor of the lower bound—exist, and are chosen at every stage. In quite a few cases, as it turns out, we find that the error term is small as compared to the rest of the denominator, and the analysis is tight if optimal or near-optimal estimators are assumed to have been used. We now show the main results in action via two examples, one where the bounds match and the error term can be ignored, and one where the error term dominates, but still leads to a gap w.r.t. the oracle-assisted loss.
IV-C Demonstrative examples for data accumulation
We first demonstrate choices for the sequence $n_{t},\alpha_{t},t\geq 0$ where the upper and lower bounds match (up to constant factors) and show a polynomially growing gap from the oracle-assisted minimax loss in the accumulation setting.
IV-C1 An example with matching bounds
Let $n_{i}=a(i+1)^{\beta+\gamma}$, and $\alpha_{i}=(i+1)^{-\gamma}$, where $a,\beta,\gamma$ are positive constants such that $a\geq\nicefrac{{k}}{{4}}$ and $\gamma\leq\min(1,\nicefrac{{1{+}\beta}}{{2}})$. Thus, on average, the number of synthetic samples at stage $t$ is $\Theta(t^{\beta+\gamma})$, while the number of real samples is $\Theta(t^{\beta})$. The assumption on $a$ ensures that the lower bound (Theorem 1) holds for every $t\geq 1$, and the assumptions on $\gamma$ ensure that the upper bound (Theorem 2) is valid and the error term $h_{\nicefrac{{1}}{{4k}}}(t)$ is, in the worst case, comparable to $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}$ for each $t$. Additionally, for the base setting, assume that the estimators in Theorem 2 are used at each stage.
Claim 1.
For the parameters described above, the minimax loss $r^{\ell_{2}^{2}}_{t,k}$ is $\Theta(t^{-(\beta-\gamma+1)})$, and the oracle-assisted minimax loss $r^{\ell_{2}^{2},\mathrm{oracle}}_{t,k}$ is $\Theta(t^{-(\beta+1)})$.
@@ IV-C1 An example with matching bounds
First, note that $\sum_{i=0}^{t}n_{i}\alpha_{i}=\sum_{i=0}^{t}a(i+1)^{\beta}=\Theta\left((t+1)^{\beta+1}\right)$. Using Lemma 5,
$\frac{c_{1}(1+\beta)}{a(t+1)^{1+\beta}}\leq r^{\ell_{2}^{2},\mathrm{oracle}}_{t,k}\leq\frac{c_{2}(1+\beta)}{a(t+1)^{1+\beta}}.$
In contrast, for the base case, using Theorem 2, since $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}=a\sum_{i=0}^{t}(i+1)^{\beta-\gamma}\geq a\frac{(t+1)^{\beta-\gamma+1}}{\beta-\gamma+1}$,
@@
Since the variance of each component $j$ of each estimate $\hat{P}_{i-1}$ is bounded above as $\text{var}\left(\hat{P}_{i-1}[j]\right)\leq E\left[\ell_{2}^{2}\left(\hat{P}_{i-1},p\right)\right]\leq r_{i-1,k}^{\ell_{2}^{2}}$ and $\hat{P}_{i-1}$ is unbiased for every $i$, using Chebyshev’s inequality on $\bar{\mathbf{P}}_{i-1,({p})}\left\{\hat{P}_{i-1}[j]<p[j]-\nicefrac{{1}}{{4k}}\right\}$,
$h_{t}(\nicefrac{{1}}{{4k}})=$ $\sum_{i=1}^{t}n_{i}\alpha_{i}g_{i}(\nicefrac{{1}}{{4k}})$
$=$ $\sum_{i=1}^{t}n_{i}\alpha_{i}\sup_{p}\max_{j}\bar{\mathbf{P}}_{i-1,({p})}\left\{\hat{P}_{i-1}[j]<p[j]-\nicefrac{{1}}{{4k}}\right\}$
$\leq$ $\sum_{i=1}^{t}a(i+1)^{\beta}\cdot\frac{\text{var}\left(\hat{P}_{i-1}\right)}{\left(\nicefrac{{1}}{{4k}}\right)^{2}}\leq\sum_{i=1}^{t}a(i+1)^{\beta}\cdot 16k^{2}\cdot r_{i-1,k}^{\ell_{2}^{2}}$
$\leq$ $\sum_{i=1}^{t}a(i+1)^{\beta}\cdot 16k^{2}\cdot\frac{\left(1-\frac{1}{k}\right)(\beta-\gamma+1)}{a\cdot i^{\beta-\gamma+1}}\leq 2^{\beta}\cdot 16k^{2}\cdot(\beta-\gamma+1)\sum_{i=1}^{t}i^{\gamma-1}$ (8)
$\leq$ $2^{\beta}\cdot 16k^{2}\cdot(\beta-\gamma+1)\frac{t^{\gamma}}{\gamma}.$
@@
This inductively shows that the upper and lower bounds match in terms of order w.r.t. the stage $t$: for each stage $t$, the upper bound is within a constant factor of the lower bound, provided that order-optimal estimators are chosen for stages 00 through $t-1$. Moreover, there is a polynomially growing gap between the base and oracle-assisted minimax losses:
$\frac{r^{\ell_{2}^{2}}_{t,k}}{r^{\ell_{2}^{2},\mathrm{oracle}}_{t,k}}=\Theta(t^{\gamma}).$ (9)
The general regime where the bounds match is given in Proposition 2.
IV-C2 An example where the upper and lower bound do not match
@@
Consider the previous example but with $1>\gamma>\frac{1{+}\beta}{2}$ instead. The upper bound on $\gamma$ ensures that the upper bound from Theorem 2 is still valid, but the error term will now dominate the analysis of the lower bound (Theorem 1). The effect of the error term growing large is seen in the following Proposition.
Claim 2.
Consider the self-consuming distribution estimation problem with the parameters given above. For $t\geq 1$, assume that, for stages $i\in[0:t-1]$, there exists a sequence of estimators for which the worst-case expected loss is smaller than $M\cdot\nicefrac{{1}}{{(i+1)^{\frac{\beta{+}1}{2}}}}$ for a positive constant $M$. If these estimators are used in the first $t$ stages, then there exists a positive constant $m$ independent of $t$ such that the minimax loss $r^{\ell_{2}^{2}}_{t,k}$ at stage $t$ is larger than $m\cdot\nicefrac{{1}}{{(t+1)^{\frac{\beta{+}1}{2}}}}$. The worst-case expected loss for the estimators in Theorem 2 remains $\mathcal{O}(\nicefrac{{1}}{{(t+1)^{\beta+1-\gamma}}})$.
Proof:
@@
Let the worst-case expected loss of the sequence of estimators at stage $t$ be denoted by $\bar{\sigma}_{t}^{2}$, with $\bar{\sigma}_{-1}^{2}=0$. Then for $t\geq 1$, using Theorem 1 and Chebyshev’s inequality (applied in the same way as (8))
$\bar{\sigma}_{t}^{2}\geq$ $\frac{1}{512a\sum_{i=0}^{t}(i+1)^{\beta-\gamma}+\frac{(i+1)^{\beta}k^{2}\bar{\sigma}_{i-1}^{2}}{32}}.$ (10)
Now, by assumption, for $i\in[0:t-1]$,
$\bar{\sigma}_{i}^{2}\leq$ $\frac{M}{(i+1)^{\nicefrac{{(\beta{+}1)}}{{2}}}}$ (11)
$\implies\sum_{i=1}^{t}(i+1)^{\beta}\bar{\sigma}_{i-1}^{2}\leq M\sum_{i=1}^{t}\left(1+\frac{1}{i}\right)^{\beta}i^{\frac{\beta-1}{2}}\leq$  $\,2^{\beta}\cdot M\sum_{i=1}^{t}i^{\frac{\beta-1}{2}}\leq 2^{\beta}\cdot M\cdot\frac{2}{\beta+1}\cdot(t+1)^{\frac{\beta+1}{2}}.$
Also, since $\beta-\gamma+1<\nicefrac{{\beta{+}1}}{{2}}$ by assumption, $\sum_{i=0}^{t}(i+1)^{\beta-\gamma}=o\left(\left(t+1\right)^{\frac{\beta+1}{2}}\right)$. Substituting these into (10),
$\bar{\sigma}_{t}^{2}\geq$ $\frac{1}{512a\cdot\frac{k^{2}}{32}\cdot 2^{\beta}\cdot M\cdot\frac{2}{\beta+1}\cdot(t+1)^{\frac{\beta+1}{2}}+o\left(\left(t+1\right)^{\frac{\beta+1}{2}}\right)},$
setting $m=512a\cdot\frac{k^{2}}{32}\cdot 2^{\beta}\cdot M\cdot\frac{2}{\beta+1}+\epsilon$, where $\epsilon$ accounts for the $o(\cdot)$ term, completes the proof.
The upper bound on the worst-case loss of the estimators in Theorem 2 follows unchanged from Example 1. ∎
Comparing the lower bound against the oracle-assisted minimax loss shows that there is a gap of at least $\Theta(t^{\nicefrac{{\beta{+}1}}{{2}}})$. The gap with the upper bound remains $\Theta(t^{\gamma})$. Since the error term $h_{t}(\nicefrac{{1}}{{4k}})$ dominates the analysis of the lower bound, the impact of a much faster decaying sequence of $\alpha_{t}$’s on the lower bound is limited. This is not to say that the impact of the penalty is small: the effective net sample size still becomes $\mathcal{O}(\sqrt{t^{\beta+1}})$ as opposed to $\Theta(t^{\beta+1})$ with oracle information. The two regimes for $\beta,\gamma\in[0,1]^{2}$ are depicted in Figure 2.
Figure 2: Regimes of minimax loss in the data accumulation workflow with $\Theta(t^{\beta})$ real and $\Theta(t^{\beta+\gamma})$ synthetic samples at stage $t$; gaps indicated are w.r.t. the oracle-assisted loss
We should also note that stronger upper bounds, for example, of the form $h_{t}(\nicefrac{{1}}{{4k}})\leq\exp(-\lambda_{0}t)$, can be assumed by restricting the class of estimators appropriately; this would then make the estimators in Theorem 2 order optimal for every choice of parameters that satisfies the hypothesis condition. Bounds using the Chebyshev inequality remain the tightest without such assumptions.
IV-D General regimes for data accumulation
We first show that when $\alpha_{t}$ is $\Omega(1)$, the minimax error coincides with the oracle-assisted minimax error up to constant factors.
Proposition 1.
When $\alpha_{t}\geq l>0$ for all $t\geq 0$, there exist constants $m$ and $M$ independent of $t$ such that
$\frac{m}{\sum_{i=0}^{t}n_{i}\alpha_{i}}\leq r^{\ell_{2}^{2}}_{t,k}\leq\frac{M}{\sum_{i=0}^{t}n_{i}\alpha_{i}}$
Proof:
It is easy to see that since $l\leq\alpha_{t}\leq 1$, $l\sum_{i=0}^{t}n_{i}\alpha_{i}\leq\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}\leq\sum_{i=0}^{t}n_{i}\alpha_{i}$. The bounds follow directly from Theorems 1 and 2. ∎
We now describe the more interesting regime where the upper and lower bounds on the minimax loss match in the following Proposition. It asserts that when $\alpha_{t}$ vanishes with $t$, the bounds match if the $\ell_{2}^{2}$ loss vanishes faster.
Proposition 2.
When $\alpha_{t}\downarrow 0$, the ratio between the upper and the lower bounds on the minimax $\ell_{2}^{2}$ loss (Theorems 2 and 1 respectively) is bounded above by a constant independent of the stage $t$ if $\alpha_{t}=\Omega\left(\nicefrac{{1}}{{\sum_{i=0}^{t-1}n_{i}\alpha_{i}^{2}}}\right)$. Moreover, if $\nicefrac{{1}}{{\sum_{i=0}^{t-1}n_{i}\alpha_{i}^{2}}}=o(\alpha_{t})$, there exists $t_{0}\geq 0$ such that the ratio is also bounded above by a constant independent of $k$ for stages $t\geq t_{0}$.
Proof:
Using Chebyshev’s inequality as shown above, along with the assumption on $\alpha_{t}$, ensures that $g_{t}(\nicefrac{{1}}{{4k}})$ is smaller than $16k^{2}\cdot m\alpha_{t}$ for a positive constant $m$, and therefore $n_{i}\alpha_{i}g_{i}(\nicefrac{{1}}{{4k}})$ is smaller than $16k^{2}\cdot mn_{i}\alpha_{i}^{2}$; the lower bound is then
$r^{l_{2}^{2}}_{t,k}\geq\frac{\nicefrac{{1}}{{512}}}{\left(1+16k^{2}m\right)\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}},$ (12)
completing the first part of the proof. If $\nicefrac{{1}}{{\sum_{i=0}^{t-1}n_{i}\alpha_{i}^{2}}}=o(\alpha_{t})$, $\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}$ also diverges (since $\alpha_{t}\leq 1$), and one can find $t_{1}$ such that $g_{t}(\nicefrac{{1}}{{4k}})$ is smaller than $\epsilon$—independent of $k$—for $t\geq t_{1}$. The denominator of (12) can then be split into $(1+16k^{2}m)\sum_{i=0}^{t_{1}-1}n_{i}\alpha_{i}^{2}$ and $(1+\epsilon)\sum_{i=t_{1}}^{t}n_{i}\alpha_{i}^{2}$, the latter of which will dominate the former as $t$ grows large. ∎
If $\alpha_{t}$ decays too quickly, the error term starts dominating the denominator of the lower bound. In spite of this, we have the following inductive lower bound on the minimax loss:
Proposition 3.
Consider the self-consuming distribution estimation problem with parameters $n_{t},\alpha_{t}$ such that
  1. 1.
    $\alpha_{t}\downarrow 0$,
  2. 2.
    $\alpha_{t}=\mathcal{O}(1/\sum_{i=0}^{t-1}n_{i}\alpha_{i}^{2})$,
  3. 3.
    $n_{t}\alpha_{t}=\mathcal{O}(\sum_{i=0}^{t-1}n_{i}\alpha_{i})$
  4. 4.
    $\sum_{i=0}^{t}n_{i}\alpha_{i}\uparrow\infty$.
Assume that there exist estimators $\hat{p}_{i}(\cdot)$ such that the worst case $\ell_{2}^{2}$ loss is bounded as
$\sup_{p}E[\ell_{2}^{2}(\hat{P}_{i},p)]\leq\frac{M}{\sqrt{\sum_{j=0}^{i}n_{j}\alpha_{j}}}$
for each $i$ in $[0:t-1]$, for some $M>0$. Then there exists a constant $m>0$ such that the minimax loss at stage $t$ is bounded below as
$r^{\ell_{2}^{2}}_{t,k}\geq\frac{m}{\sqrt{\sum_{i=0}^{t}n_{i}\alpha_{i}}}$
when the aforementioned estimators are selected for stages 00 to $t-1$.
The proof of this Proposition is deferred to Appendix D. The third condition here is satisfied by all sequences for which $n_{t}\alpha_{t}$ grows exponentially or slower (as can be seen by the Discrete Gronwall inequality [21]). This proposition solidifies the intuition on the lower bound when the real sampling probabilities $\alpha_{t}$ decay too quickly: the effective sample size is proportional to the square root of the sample size when in the oracle-assisted setting.
IV-E Data replacement
In this section, we demonstrate the application of Corollary 4 in the data replacement setting where, at every stage $t$, the estimator only has access to samples generated in the current stage $X_{t}^{n}$. Note that $\alpha$ and $n$ do not change with $t$ in this setting, except at stage 00 where all $n$ samples are real. This setting has been explored in depth in the literature e.g. [6, 9, 7, 5, 11, 10, 14, 8].
Let the worst-case expected $\ell_{2}^{2}$ loss at stage $t$ be denoted as $\bar{\sigma}_{t}^{2}$, and define $M:=\sup_{t\geq 0}\bar{\sigma}_{t}^{2}$. Repeatedly applying Corollary 4, for every $t\geq 0$,
$\bar{\sigma}_{t+1}^{2}\geq\frac{1}{512\left(n\alpha^{2}+n\alpha\sup_{p^{\prime}}\max_{j}\bar{\mathbf{P}}_{t,({p^{\prime}})}\left\{\hat{P}_{t}[j]-p^{\prime}[j]<\nicefrac{{1}}{{4k}}\right\}\right)}.$ (13)
Once again using Chebyshev’s inequality,
$\bar{\mathbf{P}}_{t,({p^{\prime}})}\left\{\hat{P}_{t}[j]-p^{\prime}[j]<\nicefrac{{1}}{{4k}}\right\}\leq\frac{E_{\bar{\mathbf{P}}_{t,({p^{\prime}})}}\left[\left(\hat{P}_{t}[j]-p^{\prime}[j]\right)^{2}\right]}{\nicefrac{{1}}{{(4k)^{2}}}}\leq 16k^{2}\cdot\bar{\sigma}_{t}^{2}\leq 16k^{2}M.$
Substituting this into (13), the smallest worst-case error has the lower bound